factor completely
20n^4-22n^3-12n^2

Answer by checkley71(8403)

20n~4-22n~3-12n~2=0 or 2n~2(10n~2-11n-6)=0 or 2n~2(5x+2)(2x-3)=0 then 2n~2-0 or n=0 & 5n+2=0 or 5n=-2 or n=-.4 & 2n-3=0 or or 2n=3 or n=3/2 or n=1.5